OpenStudy (anonymous):

sin theta=(sqrt 2)/2 Which isn't the value of theta? 225 45 135 405 All in degrees

OpenStudy (anonymous):

\[\sin \theta=\frac{ \sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Think about on the unit circle where sin(θ) has that value. It should be in two place. Obviously above the x axis because it is a positive y value. The correct answers are π/4 and 3π/4

OpenStudy (anonymous):

I think it's B 45 degrees